OpenStudy (marlee):

ax+2b=bx+c please help solve for x

OpenStudy (anonymous):

x=(c-2b) / a-b

OpenStudy (llort):

x= c-2b/a-b b=a, c=2b

jhonyy9 (jhonyy9):

ax +2b=bx+c ax-bx=c-2b x(a-b)=c-2b x=(c-2b)/(a-b)